OpenStudy (anonymous):

Can someone please help!!! I keep getting it wrong!!! Find the extreme values of f(xy)=3x^2+2y^2−4x−3 on the region described as x^2+y^2 is less than or equal to 4

OpenStudy (harsimran_hs4):

hey have you heard of langranges multipliers or not?

OpenStudy (anonymous):

yea but i still got it wrong

OpenStudy (harsimran_hs4):

ok so can you show the working ...

OpenStudy (anonymous):

yea this is what i did: f_x=6x-4<=4 6x<=8 x<=4/3 f_y=4y<=4 f_y<=1 fxx=6 fyy=4

OpenStudy (anonymous):

and then i got stuck so i assumed the max was 6 and min was 1

OpenStudy (anonymous):

but then i tried plugging in the values to the orginal equation and the number seemed wrong so idk what to do

OpenStudy (harsimran_hs4):

firstly tell me that you want to proceed using concept of maxima minima using partial derivatives of langranges multipliers

OpenStudy (anonymous):

is there an easier way

OpenStudy (anonymous):

im good with anything

OpenStudy (harsimran_hs4):

well perhaps i don`t think so because it`s gone in 2 dimension so lets do it by langranges multipliers 1. tell me what is gradient of f(x,y) ?

OpenStudy (anonymous):

fx=6x-4 fy=4y

OpenStudy (harsimran_hs4):

so that is (6x -4)i + (4y)j tell me the same for the boundary x^2 + y^2 = 4 also?

OpenStudy (anonymous):

im not sure im confused

OpenStudy (anonymous):

fx=2x and fy=2y

OpenStudy (anonymous):

so 2xi and 2yj

OpenStudy (harsimran_hs4):

yes that is fine (2x)i + (2y)j so for extreme values to occur the gradient of the function and that of the boundary must be proportional so (6x -4)i + (4y)j = c((2x)i + (2y)j) now compare the i and j components and find the value of x and y in terms of c and tell me?

OpenStudy (anonymous):

sry im not exactly sure what you mean

OpenStudy (anonymous):

and whered the c come from

OpenStudy (harsimran_hs4):

6x - 4 = 2cx and 4y = 2cy c is the proportionality constant better say the extreme values occurs when the gradient of two is antiparallel or proportional

OpenStudy (anonymous):

so 2y=c

OpenStudy (anonymous):

3x=c+4

OpenStudy (anonymous):

im sry im not sure what to do now

OpenStudy (harsimran_hs4):

x = 4/(6 - 2c) and c = 2 from second equation so x = 2 not the boundary x^2 + y^2 = 4 putting x = 2 and finding out value of y which come out to be 0 so (2, 0) is the point of extreme value

OpenStudy (anonymous):

so 2 is the max and 0 is the min

OpenStudy (anonymous):

or do we plus in 2 into the equatio and thats hte max and 0 is the min

OpenStudy (harsimran_hs4):

no the point (2, 0) will result in the extreme value of the function put x = 2 and y = 0 to yield the extreme value

OpenStudy (anonymous):

but wouldnt that result in just one value and not max and min

OpenStudy (anonymous):

yr supposed to find absolute max and min and i got the answer ot be 1 which is wrong

OpenStudy (anonymous):

@harsimran_hs4 did you get 1 as well

OpenStudy (harsimran_hs4):

yes perhaps then we screwed up with the calculation part can you tell me the answers if you have them with you?

OpenStudy (anonymous):

the answers arent given :(

OpenStudy (anonymous):

its an online hw thing

OpenStudy (harsimran_hs4):

ok then why did you say it`s wrong?

OpenStudy (anonymous):

because i plugged t in and it said it was wrong

OpenStudy (harsimran_hs4):

ok let me figure out then and i`ll get back to you

OpenStudy (anonymous):

ok thanks

OpenStudy (harsimran_hs4):

can you plug in -3 and check if thats minimum value?

OpenStudy (anonymous):

its not -3

OpenStudy (harsimran_hs4):

do you get -ve points if you put in a wrong answer if no then check out if 9 is maximum or not

OpenStudy (anonymous):

i only have one more try left to find out what the max and min is

OpenStudy (harsimran_hs4):

then wait...

OpenStudy (harsimran_hs4):

you need to plug in both the values simultaneously or not?

OpenStudy (anonymous):

yea but i can still get partial pints

OpenStudy (harsimran_hs4):

http://www.math.uri.edu/~bkaskosz/flashmo/graph3d/ check this this will give some insight to the problem but do link someone else also to this problem and figure out... i`ll try to get to some sure answer and then reply